Calculate the dead weight loss of the subsidy if the demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-13T09:07:40-0500

60-0.5p=-2+2p

-2.5p=-62

p=24.8

Q=60-12.4=47.6

Next step will need information about subsidy